OpenStudy (flvskidd):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

OpenStudy (flvskidd):

Okay I think I know the answer, I just know that my teacher will not accept the answer without the formula and steps.

OpenStudy (flvskidd):

which is confusing me

OpenStudy (anonymous):

so u just don't know the formula

OpenStudy (flvskidd):

you could say that

OpenStudy (phi):

I would use the formula rate * time = work done in this case, the work done is 1 room so rate_carl * t1 = 1 (room) and rate_jen * t2 = 1 where t1 is the time it takes carl and t2 the time for jen carl is 3 hours faster than jen which means t1 is 3 hours shorter than t2 in other words t1= t2-3

OpenStudy (flvskidd):

Here is an example similar to the question. Maybe this will ring a bell.

OpenStudy (flvskidd):

i included the steps

OpenStudy (phi):

They are using some "short cut" formula that is not easy to follow, but does work. If we ignore why it works (which is not something I usually do), and just use it, we write \[ \frac{1}{t_1}+ \frac{1}{t_2}= \frac{1}{t_{total}}\] where t1 is the time it takes for one person to do the job (working alone) t2 the amount of time for the other person to do the job working alone and t_total is the amount of time to do the job when working together

OpenStudy (flvskidd):

okay so I would plug 3 and 2 as the first two denominators?

OpenStudy (phi):

They say If they work together, they can complete the job in 2 hours that means t_total is 2 we can put that into the formula \[ \frac{1}{t_1}+ \frac{1}{t_2}= \frac{1}{t_{total}} \\ \frac{1}{t_1}+ \frac{1}{t_2}= \frac{1}{2}\]

OpenStudy (phi):

they want to find Jen's time. that is obviously unknown, so let's assume she is t1, but we will use "x" for unknown. Put x in for t1 \[ \frac{1}{x}+ \frac{1}{t_2}= \frac{1}{2} \]

OpenStudy (phi):

Now you have to figure out what this means Carl can paint a room 3 hours faster than Jennifer can what is the time t2 (Carl's time) ? we are using x for jen's time, so we can say t2 (carl's time) is 3 hours less than x

OpenStudy (phi):

do you know how to write 3 less than x in algebra?

OpenStudy (flvskidd):

3>x ?

OpenStudy (flvskidd):

3<x sorry

OpenStudy (phi):

good guess, but that is not what they mean what is 3 less than 10 ?

OpenStudy (flvskidd):

7

OpenStudy (phi):

and 3 less than 8 ?

OpenStudy (flvskidd):

5

OpenStudy (phi):

notice 3 less than 10 you figured out by doing 10-3 and for 3 less than 8 you did 8-3 what is 3 less than x?

OpenStudy (flvskidd):

-3x ?

OpenStudy (phi):

use the same idea as when you do 3 less than 5 you write 5 then a - sign then 3 5-3 =2 3 less than 5 is 2 now do 3 less than x

OpenStudy (flvskidd):

x-3

OpenStudy (phi):

yes. it is a way of "thinking" that is useful. so if jen takes x hours, carl takes x-3 hours (3 hours less than jen's x)

OpenStudy (phi):

replace t2 with x-3 in the equation \[ \frac{1}{x}+ \frac{1}{x-3}= \frac{1}{2} \]

OpenStudy (phi):

do you have to solve this ? you get a quadratic equation that you have to factor.

OpenStudy (flvskidd):

ok i understand so I solve x+x-3=2

OpenStudy (phi):

no, it's messier. one way to proceed is to multiply both sides (and all terms) by x(x-3) can you do that?

OpenStudy (flvskidd):

i think so let me try

OpenStudy (flvskidd):

sorry for being slow. if I was to multiply 2 by x(x-3) would i plug in 2 for the xs

OpenStudy (flvskidd):

sorry i dont know how to multiply each term

OpenStudy (phi):

you write x(x-3) next to each term

OpenStudy (phi):

notice you can simplify the first term because x divided by x "cancels" and in the 2nd term (x-3)/(x-3) also cancels.

OpenStudy (flvskidd):

yes I see that, would the same apply for (x-3) over (x-3)

OpenStudy (phi):

\[ \frac{x(x-3)}{x}+ \frac{x(x-3)}{(x-3)}= \frac{x(x-3)}{2} \] or after simplifying \[ (x-3) + x = \frac{x(x-3)}{2} \]

OpenStudy (flvskidd):

gotcha

OpenStudy (phi):

can you simplify the left side x-3+x

OpenStudy (flvskidd):

2x-3?

OpenStudy (phi):

yes, so you have \[ 2x-3 = \frac{x(x-3)}{2} \] if we multiply both sides by 2 we can get rid of the fraction on the right side \[ 2(2x-3) = \cancel{2}\cdot \frac{x(x-3)}{\cancel{2}} \] \[ 2(2x-3)= x(x-3) \] I would "distribute the 2" on the left side can you do that ?

OpenStudy (flvskidd):

4x-12

OpenStudy (phi):

ok except 2*3 is 6 (not 12)

OpenStudy (flvskidd):

i feel so stupid sorry

OpenStudy (phi):

\[ 4x -6 = x(x-3) \] it's easy to make simple mistakes. you have to go slower and more carefully (remember, you have to learn math and it's painful) now distribute the x on the right side

OpenStudy (flvskidd):

4x-6 = x^2 -3x

OpenStudy (phi):

now "move" the terms on the left side to the right side: add -4x+6 to both sides

OpenStudy (flvskidd):

x^2 -7x +6 ?

OpenStudy (phi):

ok, but it's an equation \[ 0= x^2 -7x +6 \] or \[ x^2 -7x +6 =0 \]

OpenStudy (flvskidd):

okay so do i factor?

OpenStudy (phi):

yes , and it does factor

OpenStudy (flvskidd):

(x-6)(x-1)

OpenStudy (phi):

yes but you should write the whole equation (x-6)(x-1)= 0 you get x=6 or x=1 x is the time it takes Jen to do the job. carl is 3 hours faster. So , even though the math has a solution x=1 (as a possibility) we can rule it out, because carl's time can't be 1-3 = -2 hours that leaves jen's time as 6 hours and carl's time as 3 hours

OpenStudy (flvskidd):

so that is the solution?

OpenStudy (phi):

yes

OpenStudy (flvskidd):

Okay thank you very much for dealing with me and helping.